# Bridging Binarization: Causal Inference with Dichotomized Continuous Exposures

### v4

@@ 1 Introduction
The ATE is not defined for continuous exposures; however, one popular method of quantifying the causal effect of a continuous exposure is by binarizing or dichotimizing the exposure variable, which enables researchers to target a parameter somewhat akin to the ATE. This method involves creating a new dummy variable that equals 1 when the exposure of interest is within some range of values and 0 otherwise. Then, researchers calculate the ATE using this new dummy variable as the exposure. We term this estimand the “binarized ATE” (BATE).
An example of this practice can be found in a 2015 paper by Hu et al., which investigated the relationship between pesticide exposure and health effects in Chinese farmers [4]. Individuals were divided into two groups: those with relatively high pesticide exposure levels (more than 50 pesticide applications in 2009–2011) and those with relatively low pesticide exposure levels (less than 50 pesticide applications in the same three year period). The researchers implemented a regression estimator including baseline covariates. They found evidence suggestive of some causal relationship between “high” versus “low” levels of pesticide use and abnormal health test results (including blood tests, nerve conduction studies, and neurological exams).
There are numerous other examples of researchers binarizing continuous exposures in public health contexts. For example, Hu et al. investigated the relationship between being in a “low-risk” group versus a “high-risk” group and risk of type 2 diabetes in women [5]. Wu et al. calculated the effect of air pollution and mortality [6]. In addition, researchers employ the binarized ATE across the social sciences, like in political science and economics. Kurtz and Lauretig investigated the effect of economic liberatlization on political protest activity [7]. Carneiro, Lokshin, and Umapathi looked at the effect of increased schooling on income in Indonesia [8].
In this paper, we show that the BATE is the same as the difference in average outcome under two related modified treatment policies (MTPs). MTPs are an example of a population-level intervention in which one considers a counterfactual world in which changes to underlying distributions of different variables are considered [9]. In particular, MTPs are policies that change the underlying distribution of the exposure variable in a population such that the new distribution depends on covariates as well as an individual’s natural value of exposure, as defined by Hanuese and Rotnitzky [10].
@@
This equivalence addresses many of the common critiques put forward by methods researchers have regarding binarization in causal inference. There have been textbook chapters written on epidemiological methods heavily dissuade researchers from categorizing/binarizing exposure variables [11]. Some researchers contend that binarization violates the Stable Unit Treatment Values Assumption (SUTVA), an assumption that some researchers make when prescribing a causal meaning to their statistical estimate of the ATE. In particular, one component of SUTVA is that there no are multiple versions of treatment [12]. VanderWeele and co-authors argue that binarization violates this assumption as encoding a continuous exposure into a binary one implies that there are several values of exposure embedded in one value of the binary exposure variable [13, 14]. Rothman, Greenland, and Lash make a similar argument, stating that such “open-ended categories” such as cut-offs are harmful because one category can contain many different exposure values or confounder effects [11]. However, the proposed MTPs circumvent these issues by defining the precise mechanism by which the distribution of the exposure changes. Specifically, one can re-write the average outcome under two seemingly deterministic binarized exposure levels in terms of the average outcomes under two policies that change the underlying distribution of exposure in a fully specified way. In addition to this critique, Bennette and Vickers contend that binarization assumes “homogeneity of risk within categories” [15]. However, the application of MTPs in this paper does not imply that the effect is the same for everyone; rather, in thinking about the effect of a modified treatment policy at a population level, we are considering a causal estimand that is an average of the exposure effect across the distribution of the exposure of interest in the target population.
@@
This type of dichotmization was discussed by Stitleman et al. [16]. These interventions were also considered by Díaz and van der Laan as one example of an intervention which imposes some type of cut-off on the distribution of exposure [17]. McCoy et al. introduced the “average regional effect,” which considers how the average outcome changes under modified treatment policies which preserve relative self-selection [18]. Rose and Shem-Tov consider binarization in the context of noncompliance [19]. Additionally, the binarized ATE was considered in the context of non-random missingness by van der Laan et al. [20]. The estimand in [20] reduces to the estimand we present when assuming missingness at random. In this paper, we aim to provide a thorough and detailed treatment of the assumptions underlying the BATE, as well as examples of when the assumptions may be more or less plausible, to provide practioners with more clarity regarding whether the BATE is the answer to their particular research question. We also present a novel estimand, the causal attributable effect of binarization, which extends the additive risk estimand considered in [16]. We argue that this estimand answers a more relevant causal question than the BATE as it compares the average outcome under one MTP to that in the observed world. In addition, we provide details for regression, inverse probability weighted, and augmented inverse probability weighted estimators that, under certain conditions outlined, will consistently estimate the BATE and the CAB. To our knowledge, these estimators and their underlying assumptions have not been discussed in the context of binarization.
@@
The rest of the paper will proceed as follows. In Section 2, we present the proof of equivalence between the binarized ATE and the difference in average outcome of modified treatment polices that restrict the value of the exposure variable while preserving relative self-selection. We also introduce a new estimand, the causal attributable effect of binarization, and discuss its relevance. Finally, we introduce regression adjustment-based, inverse probability weighting, and augmented inverse probability weighting estimators for the two estimands. Section 3 follows with a simulation study in which we implement both the regression and inverse probability weighting estimators for the binarized ATE and the new estimand. In Section 4, we provide a discussion of the results. Lastly, in Section 5, we applying the method of binarization to evaluate the effect of proximity to oil and gas wells during gestation on the probability of giving birth to an underweight newborn in the state of California.
@@ 2.1 Setup and Notation
Consider an observational study with baseline confounders $W\in\mathbb{R}^{p}$, a single exposure (possibly continuous) $A\in\mathbb{R}$, and a single outcome $Y\in\mathbb{R}$. Let the observed data be $O=(W,A,Y)$. Let supp(A)(A) and supp(W)(W) denote the support of the exposure and confounders, respectively.
@@ 2.2 Modified Treatment Policies
In causal inference, modified treatment policies (MTPs) are used to consider counterfactual worlds in which the distribution of the exposure $A$ conditional on confounders $W$ is changed in some way [10, 17]. The original measured exposure $A$ is called the “status-quo” or “observed” exposure, as it corresponds to the value of exposure an individual received in the observed world, prior to any intervention. We use $\tilde{A}$ to denote the post-intervention exposure, which can depend on $A$ and $W$. By considering changes in the underlying population distribution of the exposure, MTPs allow for stochastic interventions in addition to more traditional deterministic interventions. These types of interventions can therefore represent counterfactual worlds that are more realistic in modeling the type of behaviors that researchers may expect in a given population given a particular policy. One example of a MTP is a shift intervention, where researchers consider a policy where individuals with an observed exposure value $A=a$ would instead have a shifted exposure value $\tilde{A}=a+\delta$ for some $\delta\in\mathbb{R}$ [17]. Estimation strategies have been studied extensively for MTPs. For a comprehensive overview of modified treatment policies and nonparametric estimation strategies of causal effects, we refer the reader to Díaz et al. [22].
@@
Presume that we are analyzing the binarized data $Y,T,W$. A common practice is to regress $Y$ onto $T$ and $W$ to obtain an estimate of the regression function $E[Y|T,W]$ and then report the coefficient on $T$. This is a specific case of a more general plug-in estimation strategy where the estimand (binarized ATE; BATE) is given by
@@ 2.4 Identification and Causal Assumptions
Under these assumptions, $\Psi_{BATE}$ is equivalent to the difference in expected outcome under two modified treatment policies that affect the underlying distribution of exposure in the population. Specifically, we can consider setting $T=1$ equivalent to enacting a policy that restricts $A\in\mathcal{A}$ while preserving relative self-selection, and setting $T=0$ equivalent to enacting a policy that restricts $A\in\mathcal{A}^{\mathsf{c}}$ while preserving relative self-selection. This interpretation is completely transparent and explicitly defines an intervention, negating concerns about “different versions of treatment.” Given the (standard) identifying assumptions, this parameter is a perfectly reasonable causal contrast to use when estimating the effects of policies under which persons would have exposure probabilities in the thresholded region proportional to what they would have been without the thresholding. In the discussion, we give examples of policies where this assumption is sensible and when it is not.
@@ 2.5 New Parameter
In the previous section, we showed that the binarized ATE is equivalent to the difference between two average counterfactual outcomes under two different modified treatment policies. However, when considering the impact of a policy, it seems most natural to consider the difference between a counterfactual outcome under a MTP and the outcome under the status quo with no policy, as suggested by Hubbard and van der Laan [9]. For example, the estimand proposed by Haneuse and Rotnitzky is a difference that “quantifies the average effect in the studied population of a policy that switches each received dose…to the modified dose” [10]. As another example, when studying the effect of shift interventions, another popular MTP, researchers often estimate the difference between the average outcome under the proposed shift intervention and the observed average outcome [24, 25].
To this end, we propose a new estimand of interest, the causal attributable effect of binarizing (CAB). The CAB is akin to the causal attributable risk (CAR) of the MTPs described, generalized to non-binary outcomes. The CAR is a common estimand in public health used to examine the excess risk that results from a particular exposure [26]. It requires no data beyond what is used to calculate the binary ATE. The CAB is the difference between the expected outcome under one of the two policies and the expected outcome under the observed world:
@@
In addition, if considering implementing a particular threshold, estimating the BATE rather than the CAB could result in researchers overstating the effect of such a policy. For example, suppose a hospital is considering implementing a policy in which patients with a systolic blood pressure greater than or equal to 140 mmHg are put on a medication to reduce the blood pressure to under 140 mmHg, and doctors want to know how this would affect the overall risk of patients at the hospital experiencing cardiac arrest. Here, if researchers were to estimate and report the BATE of implementing such a threshold on blood pressure, researchers would be comparing average risk of cardiac arrests in a world where all patients has a systolic blood pressure greater than or equal to 140 mmHg to that in a world where everyone has a systolic blood pressure under 140 mmHg. It is easy to see how such an estimate could overstate the effect of implementing such a policy because the counterfactual comparison world in which everyone has blood pressure greater than or equal to 140 mmHg overstates the current disease burden. Researchers could instead calculate the CAB to compare the risk of cardiac arrest in a counterfactual world in which all patients have systolic blood pressure under 140 mmHg to that in the current world, which would more accurately reflect the reduction doctors could reasonably expect.
In addition to answering a more relevant research question, $\Psi_{CAB}$ should be convenient for researchers to consider in scenarios where they might consider $\Psi_{BATE}$. The identification assumptions are slightly weaker for the CAB than the BATE. Specifically, the positivity assumption becomes a one-way positivity assumption, which will depend on the particular causal contrast of interest. However, the two-way positivity assumption implies the one-way; therefore, any data generating process that allows for the identification of the BATE also allows for identification of the CAB. Thus, data generated by processes that satisfied conditions allowing for $\Psi_{BATE}$ to be estimated are also data in which $\Psi_{CAB}$ may be estimated. In the following section, we describe three methods by which researchers may implement estimators for $\Psi_{CAB}$, including details for calculating standard errors.
@@ 2.6 Estimators
The regression [29], IPW [30], AIPW [28], and TMLE [20] estimators are all asymptotically normal under certain conditions, allowing researchers the ability to construct valid confidence intervals based on the estimated standard errors. In Appendix A, we provide variance estimators for the proposed methods. It is important to note that, while the AIPW and TMLE estimators are doubly-robust, the influence curve based estimators presented in the Appendix for the variance presented are not doubly-robust - they are only consistent when both the outcome regression and the propensity score estimators are consistent [31]. In addition, if both the outcome regression $E[Y|T=t,W=w]$ and the propensity score $\pi_{\mathcal{A}}(w)$ are consistently estimated for $T\in\{0,1\},w\in\text{supp}(W)$, then the AIPW and TMLE estimators are efficient, meaning that they have the smallest variance as the sample size grows to infinity.
As noted above, the four estimators rely upon estimators of conditional expectations or conditional probabilities. Given that the MTPs defined above result in exposure distributions that depend on the distribution of the observed $A$, one may assume that one must first estimate the density of $A$. However, the equivalence result shows that the only relevant piece of the distribution of $A$ that researchers need to estimate $\Psi_{BATE}$ and $\Psi_{CAB}$ is the distribution of $T$, the binarized exposure. More concretely, if researchers only had access to the data $(W,T,Y)$ for a sample and no other information about the actual value of exposure $A$, they would still be able to consistently estimate $\Psi_{BATE}$ and $\Psi_{CAB}$. Our result also shows that, to estimate $\Psi_{BATE}$, one can use any standard ATE estimation technique plugging in $T$ as the binary exposure variable. However, the estimators for $\Psi_{CAB}$ take a different form, which can be found in Appendix A.
@@ 3.1 Data Generating Processes
The estimands of interest are the binarized ATE and the causal attributable effect of binarization when implementing a cut off at a exposure value of $A\geq 6$. Here, we define the CAB as $\Psi_{CAB}=E[Y_{\tilde{A}_{1}}]-E[Y]$, which contrasts the average outcome under the MTP which restricts values $A\geq 6$ and preserves relative self-selection with the average outcome in the observed world. Under the data generating process, the true parameter values are $\Psi_{BATE}=201.806$ and $\Psi_{CAB}=90.872$.
Over 5000 simulations, we drew samples of size $n=150,300,500$ from the data generating process and calculated the estimates $\widehat{\Psi}_{BATE}$ and $\widehat{\Psi}_{CAB}$ using the regression, IPW, and AIPW estimators outlined in Section 2.6. The regression estimator is consistent for this DGP because $E[Y|T,W]$ is linear in $T$ and $W$, as $W$ is binary. For a more general discussion of when the regression estimator will be consistent, see Appendix A.1. The estimate $\hat{\pi}_{\mathcal{A}}(W)$ necessary for implementing the IPW estimator was calculated using logistic regression with $W$ as the predictor and $T$ as the binary outcome, which is consistent for the propensity score for this DGP. Finally, the AIPW estimator is doubly robust. Therefore, all three estimators are consistent. We also estimated the standard errors of the estimates using the methods described in Section 2.6. Finally, we present the bias of the estimates as well as the standard error of the estimatoes over the simulations to evaluate the performance of the estimators.

### v5

@@ Abstract
Keywords: observational causal inference, continuous exposures, modified treatment policies
MSC 2020: 62D20
@@ 1 Introduction
The ATE is not defined for continuous exposures; however, one popular method of quantifying the causal effect of a continuous exposure is to binarize or dichotomize the exposure variable, which enables researchers to target a parameter somewhat akin to the ATE. This method involves creating a new dummy variable that equals 1 when the exposure of interest is within some range of values and 0 otherwise. Then, researchers calculate the ATE using this new dummy variable as the exposure. We term this estimand the “binarized ATE” (BATE).
An example of this practice can be found in a 2015 paper by Hu et al., where the authors investigated the relationship between pesticide exposure and health effects in Chinese farmers [12]. Individuals were divided into two groups: those with relatively high pesticide exposure levels (more than 50 pesticide applications in 2009–2011) and those with relatively low pesticide exposure levels (less than 50 pesticide applications in the same three year period). The researchers implemented a regression estimator including baseline covariates. They found evidence suggestive of some causal relationship between “high” versus “low” levels of pesticide use and abnormal health test results (including blood tests, nerve conduction studies, and neurological exams).
There are numerous other examples of researchers binarizing continuous exposures in public health contexts. For example, Hu et al. investigated the relationship between being in a “low-risk” group versus a “high-risk” group and risk of type 2 diabetes in women [11]. Wu et al. calculated the effect of air pollution and mortality [37]. In addition, researchers employ the binarized ATE across the social sciences, like in political science and economics. Kurtz and Lauretig investigated the effect of economic liberalization on political protest activity [15]. Carneiro, Lokshin, and Umapathi looked at the effect of increased schooling on income in Indonesia [3].
In this paper, we show that the BATE is the same as the difference in average outcomes under two related modified treatment policies (MTPs). MTPs are an example of a population-level intervention in which one considers a counterfactual world in which changes to underlying distributions of the exposure of interest are considered [13]. In particular, MTPs are policies that change the underlying distribution of the exposure variable in a population such that the new distribution depends on covariates as well as an individual’s natural value of exposure, as defined by Haneuse and Rotnitzky [9].
@@
This equivalence addresses many of the common critiques put forward by methods researchers regarding binarization in causal inference. There have been textbook chapters written on epidemiological methods heavily dissuading researchers from categorizing/binarizing exposure variables [24]. Some researchers contend that binarization violates the Stable Unit Treatment Values Assumption (SUTVA), an assumption that some researchers make when prescribing a causal meaning to their statistical estimate of the ATE. In particular, one component of SUTVA is that there are no multiple versions of treatment [25]. VanderWeele et al. argue that binarization violates this assumption as encoding a continuous exposure into a binary one implies that there are several values of exposure embedded in one value of the binary exposure variable [35, 36]. Rothman et al. make a similar argument, stating that such “open-ended categories,” like cut-offs, are harmful because one category can contain many different exposure values or confounder effects [24]. However, the proposed MTPs circumvent these issues by defining the precise mechanism by which the distribution of the exposure changes. Specifically, one can re-write the average outcome under two seemingly deterministic binarized exposure levels in terms of the average outcomes under two policies that change the underlying distribution of exposure in a fully specified way. In addition to this critique, Bennette and Vickers contend that binarization assumes “homogeneity of risk within categories” [1]. However, the application of MTPs in this paper does not imply that the effect is the same for everyone; rather, in thinking about the effect of a modified treatment policy at a population level, we are considering a causal estimand that is an average of the exposure effect across the distribution of the exposure of interest in the target population.
@@
This type of dichotomization was discussed by Stitelman et al. [30]. These interventions were also considered by Díaz and van der Laan as one example of an intervention which imposes some type of cut-off on the distribution of exposure [6]. McCoy et al. introduced the “average regional effect,” which considers how the average outcome changes under modified treatment policies which preserve relative self-selection [17]. Rose and Shem-Tov consider binarization in the context of noncompliance [23]. Additionally, the binarized ATE was considered in the context of non-random missingness by van der Laan et al. [32]. The estimand in [32] reduces to the estimand we present when assuming missingness at random. In this paper, we aim to provide a thorough and detailed treatment of the assumptions underlying the BATE, as well as examples of when the assumptions may be more or less plausible, to provide practitioners with more clarity regarding whether the BATE is the answer to their particular research question. We also present a novel estimand, the causal attributable effect of binarization (CAB), which extends the additive risk estimand considered in [30]. We argue that this estimand answers a more relevant causal question than the BATE as it compares the average outcome under one MTP to that in the observed world. In addition, we provide details for regression, inverse probability weighted, and augmented inverse probability weighted estimators that, under certain conditions outlined, will consistently estimate the BATE and the CAB. To our knowledge, these estimators and their underlying assumptions have not been discussed in the context of binarization.
@@
The rest of the paper will proceed as follows. In Section 2, we present the proof of equivalence between the BATE and the difference in average outcome of modified treatment policies that restrict the value of the exposure variable while preserving relative self-selection. We also introduce the CAB and discuss its relevance. Finally, we introduce regression adjustment-based, inverse probability weighting, and augmented inverse probability weighting estimators for the two estimands. Section 3 follows with a simulation study in which we implement the regression, inverse probability weighting, and augmented inverse probability weighting estimators for the BATE and CAB. In Section 4, we apply the method of binarization to evaluate the effect of proximity to oil and gas wells during gestation on the probability of giving birth to an underweight newborn in the state of California. Lastly, in Section 5, we provide a discussion of the results.
@@ 2.1 Setup and Notation
Consider an observational study with baseline confounders $W\in\mathbb{R}^{p}$ with $p\in\mathbb{N}$, a single exposure (possibly continuous) $A\in\mathbb{R}$, and a single outcome $Y\in\mathbb{R}$. Let the observed data be $O=(W,A,Y)$. Let supp(A)(A) and supp(W)(W) denote the support of the exposure and confounders, respectively.
@@ 2.2 Modified Treatment Policies
In causal inference, modified treatment policies (MTPs) are used to consider counterfactual worlds in which the distribution of the exposure $A$ conditional on confounders $W$ is changed in some way [9, 6]. The original measured exposure $A$ is called the “observed” or “status-quo” exposure, as it corresponds to the value of exposure an individual received in the observed world, prior to any intervention. We use $\tilde{A}$ to denote the post-intervention exposure, which can depend on $A$ and $W$. By considering changes in the underlying population distribution of the exposure, MTPs allow for stochastic interventions in addition to more traditional deterministic interventions. These types of interventions can therefore represent counterfactual worlds that are more realistic in modeling the type of behaviors that researchers may expect in a given population given a particular policy. One example of a MTP is a shift intervention, where researchers consider a policy where individuals with an observed exposure value $A=a$ would instead have a shifted exposure value $\tilde{A}=a+\delta$ for some $\delta\in\mathbb{R}$ [6]. Estimation strategies have been studied extensively for MTPs. For a comprehensive overview of modified treatment policies and nonparametric estimation strategies of causal effects, we refer the reader to Díaz et al. [5].
@@
Presume that we are analyzing the binarized data $(Y,T,W)$. A common practice is to regress $Y$ onto $T$ and $W$ to obtain an estimate of the regression function $E[Y|T,W]$ and then report the coefficient on $T$. This is a specific case of a more general plug-in estimation strategy where the estimand (binarized ATE; BATE) is given by
@@ 2.4 Identification and Causal Assumptions
Under these assumptions, $\Psi_{BATE}$ is equivalent to the difference in expected outcome under two modified treatment policies that affect the underlying distribution of exposure in the population. Specifically, we can consider setting $T=1$ equivalent to enacting a policy that restricts $A\in\mathcal{A}$ while preserving relative self-selection, and setting $T=0$ equivalent to enacting a policy that restricts $A\in\mathcal{A}^{\mathsf{c}}$ while preserving relative self-selection. This interpretation is completely transparent and explicitly defines an intervention, negating concerns about “different versions of treatment.” Given the (standard) causal identification assumptions, this parameter is a perfectly reasonable causal contrast to use when estimating the effects of policies under which persons would have exposure probabilities in the thresholded region proportional to what they would have been without the thresholding. In Section 4, we give examples of policies where this assumption is sensible and when it is not.
@@ 2.5 New Parameter
In the previous section, we showed that the BATE is equivalent to the difference between two average counterfactual outcomes under two different modified treatment policies, under causal assumptions. However, when considering the impact of a policy, it seems most natural to consider the difference between a counterfactual outcome under one MTP and the outcome under the status quo with no policy, as suggested by Hubbard and van der Laan [13]. For example, the estimand proposed by Haneuse and Rotnitzky is a difference that “quantifies the average effect in the studied population of a policy that switches each received dose…to the modified dose” [9]. As another example, when studying the effect of shift interventions, researchers often estimate the difference between the average outcome under the proposed shift intervention and the observed average outcome [18, 19].
To this end, we propose a new estimand of interest, the causal attributable effect of binarizing (CAB). The CAB is akin to the causal attributable risk (CAR) of the MTPs described, generalized to non-binary outcomes. The CAR is a common estimand in public health used to examine the excess risk that results from a particular exposure [10]. It requires no data beyond what is used to calculate the BATE. The CAB is defined as the difference between the expected outcome under one of the two policies and the expected outcome under the observed world:
@@
In addition, if considering implementing a particular threshold, estimating the BATE rather than the CAB could result in researchers overstating the effect of such a policy. For example, suppose a hospital is considering implementing a policy in which patients with a systolic blood pressure greater than or equal to 140 mmHg are put on a medication to reduce the blood pressure to under 140 mmHg, and doctors want to know how this would affect the overall risk of patients at the hospital experiencing cardiac arrest. Here, if researchers were to estimate and report the BATE of implementing such a threshold on blood pressure, researchers would be comparing average risk of cardiac arrests in a world where all patients have a systolic blood pressure greater than or equal to 140 mmHg to that in a world where all patients have a systolic blood pressure under 140 mmHg. It is easy to see how such an estimate could overstate the effect of implementing such a policy because the counterfactual comparison world in which everyone has blood pressure greater than or equal to 140 mmHg overstates the current disease burden. Researchers could instead calculate the CAB to compare the risk of cardiac arrest in a counterfactual world in which all patients have systolic blood pressure under 140 mmHg to that in the current world, which would more accurately reflect the reduction doctors could reasonably expect.
In addition to answering a more relevant research question, $\Psi_{CAB}$ should be convenient for researchers to consider in scenarios where they might consider $\Psi_{BATE}$. The identification assumptions are slightly weaker for the CAB than the BATE. Specifically, the positivity assumption becomes a one-way positivity assumption, which will depend on the particular causal contrast of interest. However, the two-way positivity assumption implies the one-way; therefore, any data generating process that allows for the identification of the BATE also allows for identification of the CAB. Thus, data generated by processes that satisfied conditions allowing for $\Psi_{BATE}$ to be estimated are also data in which $\Psi_{CAB}$ may be estimated. In the following section, we describe four methods by which researchers may implement estimators for $\Psi_{CAB}$, including details for calculating standard errors in Appendix A.
@@ 2.6 Estimators
The regression [29], IPW [16], AIPW [22], and TMLE [32] estimators are all asymptotically normal under certain conditions, allowing researchers the ability to construct valid confidence intervals based on the estimated standard errors. In Appendix A, we provide variance estimators for the proposed methods. It is important to note that, while the AIPW and TMLE estimators are doubly-robust, the influence curve based estimators presented in the Appendix for the variance presented are not doubly-robust - they are only consistent when both the outcome regression and the propensity score estimators are consistent [34]. In addition, if both the outcome regression $E[Y|T=t,W=w]$ and the propensity score $\pi_{\mathcal{A}}(w)$ are consistently estimated for $T\in\{0,1\},w\in\text{supp}(W)$, then the AIPW and TMLE estimators are semi-parametrically efficient, meaning that they have the smallest variance as the sample size grows to infinity in settings where we assume no knowledge on the distributions of the observed variables.
As noted above, the four estimators rely upon estimators of conditional expectations or conditional probabilities. Given that the MTPs defined above result in exposure distributions that depend on the distribution of the observed $A$, one may assume that one must first estimate the density of $A$. However, the equivalence result shows that the only relevant piece of the distribution of $A$ that researchers need to estimate $\Psi_{BATE}$ and $\Psi_{CAB}$ is the distribution of $T$, the binarized exposure. More concretely, if researchers only had access to the data $(W,T,Y)$ for individuals and no other information about the actual value of exposure $A$, they would still be able to consistently estimate $\Psi_{BATE}$ and $\Psi_{CAB}$. Our result also shows that, to estimate $\Psi_{BATE}$, one can use any standard ATE estimation technique plugging in $T$ as the binary exposure variable. However, the estimators for $\Psi_{CAB}$ take a different form, which can be found in Appendix A.
@@ 3.1 Data Generating Processes
The estimands of interest are the binarized ATE and the causal attributable effect of binarization when implementing a cut-off at an exposure value of $A\geq 6$. Here, we define the CAB as $\Psi_{CAB}=E[Y_{\tilde{A}_{1}}]-E[Y]$, which contrasts the average outcome under the MTP which restricts values $A\geq 6$ and preserves relative self-selection with the average outcome in the observed world. Under the data generating process, the true parameter values are $\Psi_{BATE}=201.806$ and $\Psi_{CAB}=90.872$.
Over 5000 simulations, we drew samples of size $n=150,300,500$ from the data generating process and calculated the estimates $\widehat{\Psi}_{BATE}$ and $\widehat{\Psi}_{CAB}$ using the regression, IPW, and AIPW estimators discussed detailed in Appendix A. The regression estimator is consistent for this DGP because $E[Y|T,W]$ is linear in $T$ and $W$, as $W$ is binary. For a more general discussion of when the regression estimator will be consistent, see Appendix A.1. The estimate $\hat{\pi}_{\mathcal{A}}(W)$ necessary for implementing the IPW estimator was calculated using logistic regression with $W$ as the predictor and $T$ as the binary outcome, which is consistent for the propensity score for this DGP. Finally, the AIPW estimator is doubly robust. Therefore, all three estimators are consistent. We also estimated the standard errors of the estimates using the methods described in Section 2.6. Finally, we present the bias of the estimates as well as the standard error of the estimators over the simulations to evaluate the performance of the estimators.